These were accepted.
US have a list, which Alison will circulate. There is also a wiki page for parked kit which Alastair is happy to have people add things to.
Angus has been away. We think he's back now.
Please remember to update the corresponding rack diagrams:
New action on Alastair to chase this up with Dave B (see CCPAG report below)
(see CCPAG report below)
Blog articles suggested:
Report from Computing Executive Group
College will be adopting these and putting them on their wiki.
Brian Gilmore has rewritten his document. There are still some questions over: School reporting models; incident reporting; pentration testing.
Some Schools had seen consumables arriving late.
There was some doubt as to whether the "cloud printing" service was ready yet for prime time.
There probably isn't now going to be a lcfg component produced elsewhere in the College that we could use. We should try to keep the meetings going, however.
IS are looking at the possibility of an institutional drupal service. There'll be a meeting soon.
Maths and Biology both expressed an interest. Discussion was deferred.
Brian McGregor is doing a review of how this went. There were apparently quite a few complaints about EUCLID. We (at the Operational meeting) hadn't seen this, but don't interact with EUCLID in quite the same way as DoSses and the ITO.
Dave Berry hadn't heard of this issue. Other Schools will definitely need this to be available too. Alastair will continue to chase up with Dave B.
Tony is going to organise a "network capacity" meeting.
Network Services are taking over management of the phones system.
See recent email from Fiona Littleton to the ITPF.
Reports from units
This is an interim solution. Stephen, Neil and Toby will discuss how to do the package management.
The problem whereby DNS notifies to the lcfg slaves were being ignored because the source IP address used by the master didn't match what they were expecting has been fixed. The master now explicitly sets the source address to DNSmaster.net.inf.ed.ac.uk to match the "master" statements in the slaves' named.conf files.
Watch out for funnies with this driver. It'll be replaced soon, though it'll take a reboot to swap the kernel modules over.
Graham will package up his script and update the wiki page to recommend using the buildhosts.
There was much discussion as to the possible implications of this failing yet again, and a recommendation that the Services unit should implement any fixes (firmware upgrades, card replacements) as soon as possible. Neil to arrange this.
The question was raised of whether there were any circular dependencies between the AFS service, which would have to be taken down, and the package update service. There should be workarounds if so. Neil will discuss with Stephen prior to the upgrade.
Topics for discussion
It'll be on Tuesday 11th December at 13:00. Stephen will be booking shortly.
Graham gave a quick report of his discussions with the environmental consultant who is looking at the building's ventilation issues.
Perdita's recent email had covered most of the salient points in a general way. CEG will consider specifics. We suspect we're unlikely to get into a similar situation again any time soon.
The next meeting will be on Wednesday 28th November at 10:00 in IF-4.31.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh